OpenStudy (anonymous):

help please...A scientist wants to make a solution of tribasic sodium phosphate, Na3PO4, for a laboratory experiment. How many grams of Na3PO4 will be needed to produce 550mL of a solution that has a concentration of Na+ ions of 0.500M ?

OpenStudy (festinger):

Calculate the Mr of Na3PO4 Calculate the amount of mols of Na3PO4 needed. This could be tricky so I will list the steps for you check 1Na3PO4 -> 3Na+ 0.5M = 0.5mol/L So the number of mols of Na+ is: \[0.5\frac{mol}{L} * \frac{L}{1000mL}*550mL = 0.275mol\] But \[\frac{mol\, of\, Na_{3}PO_{4}}{mol\, of\, Na^+} = \frac{1}{3}\]So \[Mol\, of\, Na_{3}PO_{4} = {mol\, of\, Na^+}*\frac{mol\, of\, Na_{3}PO_{4}}{mol\, of\, Na^+} \] Which is \[Mol\, of\, Na_{3}PO_{4} = \frac{1}{3} * 0.275mol\] With this in hand you should be able to obtain the mass by using the relation \[M_{r} * mol = mass\]

OpenStudy (anonymous):

thank you very muc